# Charting the Landscape of African NLP: Mapping Progress and Shaping the Road Ahead

### v2

@@ Appendix D Other low-resource techniques
Annotation Projection (AP) which is a technique used to transfer labels from a labeled source dataset to an unlabeled target dataset is also used for African languages. Text classification benchmarks covering several African languages like SIB-200 (Adelani et al., 2024) and Taxi-1500 (Ma et al., 2023) are based on this technique, and it has also been shown to work for sequence labeling (García-Ferrero et al., 2023).
Ensemble Learning a technique which multiple models usually either by aggregating predictions or representations—is also used in some of the surveyed literature, including approaches like model fusion (Rathore et al., 2023; Tran et al., 2021).
@@
Active Learning a machine learning approach where models dynamically selects the data to train was shown to be effective for African languages in tasks such as morphophonological processing (Mirbostani et al., 2023), NER (Kholodna et al., 2024), language modeling (Dossou et al., 2022), and MT (Chimoto and Bassett, 2022).
@@
Dictionary-based NLP There are a few papers in this category. Some focus on creating lexicons for African languages Emezue et al. (2024), while others explore how to use these existing lexicons—particularly for tasks such as data augmentation (Reid et al., 2021; Wang et al., 2022). Lexicon-based augmentation has proven useful for generating more data for language modeling. Additionally, a few works have shown their usefulness in tasks such as sentiment and emotional analysis (Teodorescu and Mohammad, 2023; Mabokela et al., 2023) for African languages.
@@
Appendix E Comprehensive list of African language resources available
@@ E.1 Comprehensive list of datasets and benchmarks
In Table 3, we provide a list of some of the created and publicly released labeled and unlabeled text and audio datasets for African languages across various tasks. While some of these datasets are part of large, curated multilingual language resources, others are tailored to a single language or a small group of languages. The list also includes datasets created for African variants of English (Afonja et al., 2021; Eisenstein et al., 2023; Ogun et al., 2024a). Although a few works were done on African French (Alumäe et al., 2023; Maison and Estève, 2023) and Portuguese (Hagemeijer et al., 2022b), there were no datasets created.
@@
Although they are not highlighted in the tables, there are also multimodal, multilingual, and multitask benchmarks that have been created over the years surveyed—especially in the era of LLMs—to evaluate their generalization and reasoning abilities. They include benchmarks such as BLEnD (Myung et al., 2024b), XTREME (Hu et al., 2020), XTREME-UP (Ruder et al., 2023), IrokoBench (Adelani et al., 2025), MEGAVERSE (Ahuja et al., 2024), and AfroBench (Ojo et al., 2023). There are those specifically tailored towards dialects of languages includes dialects of African langauges (Mousi et al., 2024; Faisal et al., 2024) and also creoles which are spoken across Africa (Robinson et al., 2024; Lent et al., 2024).
@@
Aside dataset creation, there are also papers focused on data curation methodology of African languages, identifying challenges and proposing solutions such as creating platform to support creation and storage Tchiaze et al. (2020); Griesel and Bosch (2020); Griscom (2020); Marivate et al. (2020).
@@
Amharic TC (Azime and Mohammed, 2021) text classification 1
@@
MasakhaNER 2.0 Adelani et al. (2022b) IE 20
MasakhaNER Adelani et al. (2021a) IE 10
@@
MIRACL Zhang et al. (2023) IR 2
Mr. TyDi Zhang et al. (2021) IR 1 (swh)
@@
CrossSum Bhattacharjee et al. (2023) Summarization 6
XL-Sum Hasan et al. (2021) Summarization 6
@@
FLEURS Conneau et al. (2022) ASR 20
FLEURS-R Ma et al. (2024) TTS 20
SpeechTaxi Keller and Glavas (2024) Speech Classification 8
@@
Nicolingua Doumbouya et al. (2021) ASR 3
Kallaama Gauthier et al. (2024) ASR 3
@@
CulturaX LM 7
@@
MMS ulab v2 Pratap et al. (2024); Chen et al. (2024c) SRL 1450
@@
Masked Language Models
@@
Walia-LLM (Tonja et al., 2024b) Decoder LM 1 (amh)
@@
E.2 Comprehensive list of models, toolkits and platforms
Several models have been developed and released for African languages in the period surveyed, spanning a range of NLP tasks. Two key areas that have seen a proliferation of models are machine translation and language modeling. These models include both language-specific (Emezue and Dossou, 2020b; Adelani et al., 2022a) and multilingual systems (Fan et al., 2020; Elmadany et al., 2024; Mohammadshahi et al., 2022a), largely enabled by the creation of extensive multilingual datasets—such as web-mined bitexts.
While these multilingual resources can be used directly, multiple studies have shown that such models tend to underperform on African languages compared to high-resource languages such as German, French, and English (Adelani et al., 2024, 2025). To address this disparity, transfer learning has emerged as an effective strategy: pretrained multilingual models are adapted to low-resource African languages via task-specific fine-tuning or adaptive pretraining, followed by fine-tuning. These transfer learning techniques help bridge the performance gap by allowing models to leverage knowledge from high-resource languages or related tasks.
Examples of pretrained translation models used in this way include M2M-100 (Fan et al., 2020) and NLLB (NLLB Team et al., 2024), which support a wide array of African languages. For language modeling, models vary across architectural paradigms—encoder-only (Ogueji et al., 2021; Oladipo, 2024; Conneau et al., 2020b; Imani et al., 2023), decoder-only (Xue et al., 2021; Adebara et al., 2024), and encoder-decoder—and in their scope as either monolingual (Abdaoui et al., 2021; Chukwuneke et al., 2022; Martin et al., 2022; Nzeyimana and Niyongabo Rubungo, 2022; Olawole et al., 2024) or multilingual systems (Ogueji et al., 2021; Oladipo, 2024; Adebara et al., 2023b).
In recent years—including the period covered by this survey—there has been a surge in the development of general-purpose large language models (LLMs), often multilingual and containing billions of parameters. These models support zero-shot and few-shot learning through the in-context learning paradigm, making them adaptable to a wide range of downstream tasks with minimal task-specific supervision.
These LLMs include both open-source and proprietary models. While the training data and language coverage for proprietary models are often undisclosed, open models are typically trained on data from high-resource languages. Evaluations of both open and closed models on a range of tasks in African languages consistently show relatively poor performance, highlighting persistent challenges in cross-lingual generalization and data representation for low-resource languages.
We highlight several models names in LABEL:{tab:app_resource2}. For models and datasets covering a single language, we include the ISO-3 code for that language. In cases where the data pertains to dialects of non-indigenous African languages, such as African Portuguese (APs) and African English (AEAs), we use appropriate dialectal labels instead.

### v3

@@ Appendix D Other low-resource techniques
Meta Learning has increasingly been applied to African and other low-resource languages, addressing challenges like scarce labeled data (Lux et al., 2024), noisy supervision (Zhu et al., 2023), and cross-lingual transfer (Wu et al., 2023). Approaches range from cross-lingual adaptation in NLP tasks (e.g., text classification (Zhu et al., 2023), question answering (Wu et al., 2023), dependency parsing (Choenni et al., 2023), text generation (Maurya and Desarkar, 2022)) to low-resource speech recognition (Chen et al., 2024d, e) and even text-to-speech synthesis (Lux et al., 2024), using techniques such as MAML, task-based meta-loss optimization, adversarial representation learning, and dynamic subnetworks. These methods enable models to generalize from resource-rich languages to underrepresented African languages, improve performance in zero- or few-shot settings, and handle noisy or weak supervision effectively.
Annotation Projection (AP) a technique used to transfer labels from a labeled source dataset to an unlabeled target dataset is also used for African languages. Text classification benchmarks covering several African languages like SIB-200 (Adelani et al., 2024) and Taxi-1500 (Ma et al., 2023) are based on this technique, and it has also been shown to work for sequence labeling (García-Ferrero et al., 2023).
Ensemble Learning a technique that combines multiple models usually either by aggregating predictions or representations—is also used in some of the surveyed literature, including approaches like model fusion (Rathore et al., 2023; Tran et al., 2021).
@@
Active Learning a machine learning approach where models dynamically select the data to train-was shown to be effective for African languages in tasks such as morphophonological processing (Mirbostani et al., 2023), NER (Kholodna et al., 2024), language modeling (Dossou et al., 2022), and MT (Chimoto and Bassett, 2022).
@@
Dictionary-based NLP There are a few papers in this category. Some focus on creating lexicons for African languages (Emezue et al., 2024), while others explore how to use these existing lexicons—particularly for tasks such as data augmentation (Reid et al., 2021; Wang et al., 2022). Lexicon-based augmentation has proven useful for generating more data for language modeling. Additionally, a few works have shown their usefulness in tasks such as sentiment and emotional analysis (Teodorescu and Mohammad, 2023; Mabokela et al., 2023) for African languages.
@@
Appendix E Comprehensive list of available African language resources
@@ E.1 Comprehensive list of datasets and benchmarks
In Table˜3, we provide a list of some of the created and publicly released labeled and unlabeled text and audio datasets for African languages across various tasks. While some of these datasets are part of large, curated multilingual language resources, others are tailored to a single language or a small group of languages. The list also includes datasets created for African variants of English (Afonja et al., 2021; Eisenstein et al., 2023; Ogun et al., 2024a). Although a few works were done on African French (Alumäe et al., 2023; Maison and Estève, 2023) and Portuguese (Hagemeijer et al., 2022b), there were no datasets created.
@@
Although they are not highlighted in the tables, there are also multimodal, multilingual, and multitask benchmarks that have been created over the years surveyed—especially in the era of LLMs—to evaluate their generalization and reasoning abilities. They include benchmarks such as BLEnD (Myung et al., 2024b), XTREME (Hu et al., 2020), XTREME-UP (Ruder et al., 2023), IrokoBench (Adelani et al., 2025), MEGAVERSE (Ahuja et al., 2024), and AfroBench (Ojo et al., 2025). There are those specifically tailored towards dialects of languages includes dialects of African languages (Mousi et al., 2024; Faisal et al., 2024) and also creoles which are spoken across Africa (Robinson et al., 2024; Lent et al., 2024).
@@
In addition to data collection, there are also articles focused on African language data curation methodology, identifying challenges and proposing solutions such as creating platforms to support creation and storage (Tchiaze et al., 2020; Griesel and Bosch, 2020; Griscom, 2020; Marivate et al., 2020).
E.2 Comprehensive list of models, toolkits and platforms
Several models have been developed and published for African languages in the surveyed period, spanning a variety of NLP tasks. Two key areas that have seen a proliferation of models are machine translation and language modeling. These models include both language-specific (Emezue and Dossou, 2020b; Adelani et al., 2022a) and multilingual systems (Fan et al., 2020; Elmadany et al., 2024; Mohammadshahi et al., 2022a), largely enabled by the creation of extensive multilingual datasets—such as web-mined bi-texts.
Although these multilingual resources can be used directly, multiple studies have shown that such models tend to underperform in African languages compared to high-resource languages such as German, French and English (Adelani et al., 2024, 2025). To address this disparity, transfer learning has emerged as an effective strategy: pretrained multilingual models are adapted to low-resource African languages via task-specific fine-tuning or adaptive pretraining, followed by fine-tuning. These transfer learning techniques help bridge the performance gap by allowing models to leverage knowledge from high-resource languages or related tasks.
Examples of pretrained translation models used in this way include M2M-100 (Fan et al., 2020) and NLLB (NLLB Team et al., 2024), which support a wide range of African languages. For language modeling, models vary across architectural paradigms—encoder-only (Ogueji et al., 2021; Oladipo, 2024; Conneau et al., 2020b; Imani et al., 2023), decoder-only (Xue et al., 2021; Adebara et al., 2024), and encoder-decoder—and in their scope as monolingual (Abdaoui et al., 2021; Chukwuneke et al., 2022; Martin et al., 2022; Nzeyimana and Niyongabo Rubungo, 2022; Olawole et al., 2024) or multilingual systems (Ogueji et al., 2021; Oladipo, 2024; Adebara et al., 2023b).
In recent years, including the period covered by this survey, there has been a surge in the development of general-purpose LLMs, often multilingual and containing billions of parameters. These models support zero-shot and few-shot learning through the in-context learning paradigm, making them adaptable to a wide range of downstream tasks with minimal task-specific supervision.
These LLMs include open-source and proprietary models. Although the training data and language coverage for proprietary models are often undisclosed, open models are typically trained on data from high-resource languages. Evaluations of both open and closed models on a range of tasks in African languages consistently show relatively poor performance, highlighting persistent challenges in cross-lingual generalization and data representation for low-resource languages.
We highlight several models’ names in LABEL:{tab:app_resource2}. For models and datasets covering a single language, we include the ISO-3 code for that language. In cases where the data pertain to dialects of non-indigenous African languages, such as African Portuguese (APs) and African English (AEAs), we use appropriate dialectal labels instead.
@@ E.2 Comprehensive list of models, toolkits and platforms
Amharic TC (Azime and Mohammed, 2021) text classification 1 (amh)
Emakhuwa TC (Ali et al., 2024a) text classification 1 (vmw)
@@
MasakhaNER 2.0 (Adelani et al., 2022b) IE 20
MasakhaNER (Adelani et al., 2021a) IE 10
@@
MIRACL (Zhang et al., 2023) IR 2
Mr. TyDi (Zhang et al., 2021) IR 1 (swh)
@@
CrossSum (Bhattacharjee et al., 2023) Summarization 6
XL-Sum (Hasan et al., 2021) Summarization 6
@@
FLEURS (Conneau et al., 2022) ASR 20
FLEURS-R (Ma et al., 2024) TTS 20
SpeechTaxi (Keller and Glavas, 2024) Speech Classification 8
@@
Nicolingua (Doumbouya et al., 2021) ASR 3
Kallaama (Gauthier et al., 2024) ASR 3
@@
CulturaX (Nguyen et al., 2024b) LM 7
CulturaY (Thuat Nguyen and Nguyen, 2024) LM 3
HPLT (de Gibert et al., 2024) LM 3
TLMD (Gaim et al., 2021) LM 1 (tir)
@@
MMS ulab v2 (Pratap et al., 2024; Chen et al., 2024c) SRL 1450
@@
Encoder Language Models
@@
LLaMAX (Lu et al., 2024) Decoder LM 20
LLaMAX-Alpaca (Lu et al., 2024) Decoder LM/MT 20
AfriInstruct (Uemura et al., 2024) Decoder LM 19
@@
Walia-LLM (Azime et al., 2024) Decoder LM 1 (amh)